# DriveAction: A Benchmark for Exploring Human-like Driving Decisions in VLA Models

### v1

@@ Abstract
Vision-Language-Action (VLA) models have advanced autonomous driving, but existing benchmarks still lack scenario diversity, reliable action-level annotation, and evaluation protocols aligned with human preferences. To address these limitations, we introduce DriveAction, the first action-driven benchmark specifically designed for VLA models, comprising 16,185 QA pairs generated from 2,610 driving scenarios. DriveAction leverages real-world driving data proactively collected by users of production-level autonomous vehicles to ensure broad and representative scenario coverage, offers high-level discrete action labels collected directly from users’ actual driving operations, and implements an action-rooted tree-structured evaluation framework that explicitly links vision, language, and action tasks, supporting both comprehensive and task-specific assessment. Our experiments demonstrate that state-of-the-art vision-language models (VLMs) require both vision and language guidance for accurate action prediction: on average, accuracy drops by 3.3% without vision input, by 4.1% without language input, and by 8.0% without either. Our evaluation supports precise identification of model bottlenecks with robust and consistent results, thus providing new insights and a rigorous foundation for advancing human-like decisions in autonomous driving.
@@ 1 Introduction
Early autonomous driving systems were predominantly designed with a modular architecture [1, 2, 3, 4, 5, 6, 7], separating perception, prediction, planning, and control into independently optimized components. Recently, advances in large-scale multi-modal data and computational power have led to new paradigms, including end-to-end approaches and Vision-Language-Action (VLA) models [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], significantly enhancing system generalization and complex task performance. Despite these advancements, current systems still struggle with real-world diversity and replicating human driving preferences. Accordingly, within the VLA paradigm, comprehensive and rigorous evaluation of the entire pipeline is increasingly crucial in both academia and industry.
@@
Recent benchmarks and datasets[22, 25, 26, 27, 28, 29, 30] represent significant progress, yet they still struggle to capture the diversity, complexity, and behavioral characteristics of real-world driving. Most existing benchmarks [25, 26, 27, 28] are constructed from open-source datasets [31, 32, 33, 34], resulting in limited source variety. Typically, these datasets are designed for object-level perception tasks and thus overlook the contextual richness and human intent intrinsic to realistic driving decisions. In addition, critical and challenging scenarios—such as road merges and exits, interactions with pedestrians, and construction zones—remain largely underrepresented, making evaluation results less relevant to practical deployment risks. Furthermore, the distribution of user behaviors is highly imbalanced, with simple maneuvers like going straight dominating the data, while more complex events are insufficiently covered, leading to inadequate assessment of challenging behaviors.
@@
Existing approaches to action ground truth annotation exhibit several limitations. Some works [25, 26] do not provide action-level annotations and focus only on perception or understanding tasks. Other approaches [28, 35, 36, 37], utilize manually annotated action labels, but such labels are often generated after driving behavior occurs and therefore do not faithfully reflect real-time driving intent and decisions. This gap in high-fidelity action labels restricts the reliability and realism of current evaluation.
@@
Regarding the design of evaluation systems, most existing benchmarks do not fully capture the core logic of driving decision-making. Some works focus primarily on isolated tasks such as object recognition [25], video captioning [35], or spatial understanding [26], without systematically covering the entire process from vision to action. For benchmarks that do address the full decision pipeline [28, 37], a forward logic is often adopted, starting from perception and proceeding sequentially through prediction, planning, and action modules. However, this may not adequately represent a goal-driven paradigm that considers dependencies from the perspective of final decisions. As a result, current evaluation standards may not be closely aligned with realistic human driving decisions, highlighting the need for more comprehensive and goal-oriented evaluation frameworks.
@@
  • User-Contributed Broad-Coverage Driving Scenarios. DriveAction is constructed from real-world data proactively collected by internal test users of production-level autonomous vehicles, which fundamentally distinguishes it from existing benchmarks and provides a wide spectrum of both everyday and challenging driving scenarios. Manual curation guarantees a comprehensive and representative collection of driving scenarios and actions.
  • Human Driving Preference-Aligned Ground Truth. Action labels are collected directly from users’ real-time driving operations, faithfully capturing human intent at the moment of decision-making. To align with the output granularity of end-to-end large models, these labels are discretized into high-level actions, which reflect the categorical nature of human driving decisions. All labels are manually verified to ensure validity, with erroneous, unreasonable, or illegal behaviors excluded.
@@ 2.1 Autonomous Driving Models
Autonomous driving was initially approached through modular systems [1, 2, 3, 4, 5, 6, 7], in which perception, prediction, planning, and control modules were developed and optimized independently. Subsequently, research progressed towards end-to-end architectures [8, 9, 10, 11, 12, 13, 14, 15, 16] that directly learn mappings from raw sensory inputs to actions. With advances in vision-language models (VLMs)[38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59], hybrid approaches have emerged, in which VLMs are integrated into end-to-end architectures as independent modules that provide low-frequency driving suggestions [17]. Most recently, the VLA paradigm [18, 19, 20], has established, enabling deeper integration of end-to-end models and VLMs for richer context understanding and greater generalization, further highlighting the growing need for dedicated and comprehensive evaluation benchmarks.
@@ 2.2 Language-Related Benchmarks for Autonomous Driving
Existing works can be grouped into three categories. The first focuses on video captioning and behavior explanation, represented by BDD-X [35], which links actions with textual descriptions, and DRAMA [36], which uses question-answer annotations to identify risk objects and their corresponding causes. The second emphasizes 3D perception and spatial understanding. For example, NuScenesQA [25] is designed for object-level and multi-modal question answering, while DriveMLLM [26] evaluates spatial and localization capabilities. The third focuses on evaluation across the entire autonomous driving pipeline, including Reason2Drive [27], DriveLM [28], and DriveBench [37], which use a forward logic starting from perception, thus insufficiently aligned with goal-driven dependencies and realistic human driving behaviors.
@@
BDD-X [35] Caption/explanation Self-collected Manual None
DRAMA [36] Caption/explanation Self-collected Manual Chain
NuScenesQA [25] 3D perception/spatial nuScenes None None
DriveMLLM [26] 3D perception/spatial nuScenes None None
Reason2Drive [27] AD System nuScenes, Waymo, ONCE Open source Chain
DriveLM [28] AD System nuScenes, CARLA Manual Graph
DriveBench [37] AD System DriveLM Manual Graph
DriveAction AD System Production vehicle user-contributed Real-time operations Tree
@@ 3 DriveAction
Inspired by existing benchmarks, we introduce DriveAction, the first action-driven benchmark specifically designed for VLA models, leveraging real-world user driving preferences. The following sections provide a comprehensive description of DriveAction. Specifically, Section 3.1 highlights how user-contributed and carefully curated data enable extensive scenario coverage and diverse action representation. Section 3.2 details action labels aligned with human-like decision making, collected from real-time driving operations and validated through multi-stage review. Section 3.3 introduces the action-rooted, tree-structured evaluation framework, supporting flexible and rigorous assessment of models across the full V-L-A pipeline.
@@ 3.1 Driver-Contributed Broad-Coverage Driving Scenarios
As shown in the Source column of Table 1, DriveAction uniquely aggregates real-world data collected by users of autonomous driving systems in mass-produced vehicles, in contrast to previous benchmarks that rely on self-collected or open-source data. Our dataset covers 148 cities throughout China and includes records of the complete lineup of mass-produced vehicles in our deployment. To ensure a comprehensive and representative collection of driving scenarios and actions, we perform multiple rounds of manual selection and quality control.
Table 2 summarizes the seven key scenario categories in DriveAction, each paired with representative actions and concise descriptions to illustrate the coverage and annotation strategy. These scenarios span a wide range of real-world driving conditions, including ramp and side road merging/splitting, as well as navigation- and efficiency-driven lane changes. The coverage extends beyond typical urban, highway, and mixed-traffic environments to encompass challenging contexts such as complex intersections, construction zones, congestion, and interactions with vulnerable road users. Each scenario is associated with a variety of fine-grained actions—such as lane changes, deceleration, and bypass maneuvers—enabling detailed analysis of decision-making across diverse driving situations.
@@
Bypass VRU Bypass left/right For bypassing vulnerable road users (e.g., pedestrians and cyclists)
@@ 3.2 Human Driving Preference-Aligned Ground Truth
DriveAction derives action labels directly from real user driving operations, enabling accurate, real-time capture of driver intent and decisions. This contrasts with previous benchmarks, which rely on manual or open-source annotation, as indicated in the Label column of Table 1.
DriveAction adopts discretized high-level actions as ground truth, which matches the output granularity of end-to-end large models and better reflects the diversity and categorical nature of human driving decisions. For example, in lane change scenarios, instead of relying on dense trajectories sampled at high frequency, DriveAction captures decisions made at key points—such as whether to initiate a lane change and which direction to take—better matching the lower decision frequency of large models. This design provides a more suitable and fair evaluation standard for current autonomous driving models.
@@ 3.3 Action-Rooted Tree-Structured Evaluation
To establish an evaluation framework that closely mirrors real-world driving decision processes and systematically assesses autonomous driving models across the entire decision-making pipeline, we propose an action-rooted, tree-structured evaluation architecture in DriveAction. By adopting an action-rooted hierarchical design anchored in action decisions, this framework dynamically maps complex driving actions to the required vision and language tasks. Through the integration of rich contextual scenario information, it ensures that model decisions are made within a complete and realistic environment. Furthermore, the evaluation system is designed to be highly flexible, supporting both comprehensive V-L-A assessment and task-specific evaluation, and thus accommodates a wide range of model capabilities and evaluation needs.
@@ 3.3.1 Task Definition
Figure 1 illustrates the tree-structured task architecture of DriveAction, which uses the action space as the root, explicitly defines the corresponding language tasks required for each action, and further maps each language task to its dependent vision tasks. This hierarchy is organized into three layers: the top layer consists of action nodes, such as lane change and intersection turning, representing final decision outputs of the model. The middle layer consists of language tasks, such as navigation following and traffic light following, which provide scene understanding for each action. The bottom layer comprises vision tasks, responsible for detecting and recognizing key environmental elements, such as lanes, traffic signs, and obstacles. Although the evaluation framework is modeled as a dependency tree descending from actions, the underlying model inference still follows the conventional V-L-A order. This design offers a systematic and targeted evaluation structure, thus closely matching the information flow and reasoning processes of real-world autonomous driving systems.
DriveAction comprises 14 independent tasks, including 7 vision tasks and 7 language tasks, covering diverse scenarios encountered in real-world driving. All tasks are evaluated in a question-answering (QA) format with paired questions and answers, including both selection and judgment modes. Figure 2 presents the QA distribution across vision, language, and action levels, ensuring sufficient and representative coverage within each layer of the VLA hierarchy. For instance, turn actions constitute a significant portion, which aligns with their relatively high frequency in real-world driving scenarios. This distribution ensures that the evaluation covers both common and less frequent situations, providing a meaningful basis for assessing model performance. All task definitions and example QA pairs are detailed in the Appendix.
@@
Figure 3 uses the traffic sign task as an example to illustrate the complete V-L-A evaluation pipeline in DriveAction. The process begins with the detection of traffic signs, where the model is required to identify both the presence and type of each sign. Upon successful detection, the model must then interpret whether the sign is relevant to the current driving context and its own behavior. If the traffic sign is determined to have an actual impact, the model is expected to make an appropriate driving decision accordingly. Through this V-L-A pipeline, DriveAction comprehensively evaluates the model’s ability to transition from low-level perception to high-level understanding, and ultimately to informed decision-making.
@@ 3.3.2 Scenario Information Design
DriveAction incorporates key scenario context into each evaluation prompt, ensuring that models are assessed under realistic and representative conditions, and mitigating unsupported or hallucinatory reasoning. Specifically, three types of scenario information are provided:
@@
  • Vehicle Speeds: Ego and target vehicle speeds, obtained from onboard sensors, quantify both the current and desired driving states. This dynamic information, which cannot be inferred from images alone, is essential for making rational decisions in lane changes, overtaking, or acceleration.
To evaluate the effect of scenario information, we present a lane change task before an intersection under two conditions: with and without navigation instructions. As shown in Figure 4, the model accurately selects the correct lane change action in advance when the navigation instruction clearly indicates an upcoming right turn, while in the absence of this information, it often makes "hallucinated" decisions that do not align with the actual driving goal. This highlights the importance of providing key scenario information for reliable and context-aware autonomous driving evaluation.

### v2

@@ Abstract
Vision-Language-Action (VLA) models have advanced autonomous driving, but existing benchmarks still lack scenario diversity, reliable action-level annotation, and evaluation protocols aligned with human preferences. To address these limitations, we introduce DriveAction, the first action-driven benchmark specifically designed for VLA models, comprising 16,185 QA pairs generated from 2,610 driving scenarios. DriveAction leverages real-world driving data proactively collected by drivers of autonomous vehicles to ensure broad and representative scenario coverage, offers high-level discrete action labels collected directly from drivers’ actual driving operations, and implements an action-rooted tree-structured evaluation framework that explicitly links vision, language, and action tasks, supporting both comprehensive and task-specific assessment. Our experiments demonstrate that state-of-the-art vision-language models (VLMs) require both vision and language guidance for accurate action prediction: on average, accuracy drops by 3.3% without vision input, by 4.1% without language input, and by 8.0% without either. Our evaluation supports precise identification of model bottlenecks with robust and consistent results, thus providing new insights and a rigorous foundation for advancing human-like decisions in autonomous driving.
@@ 1 Introduction
Early autonomous driving systems were predominantly designed with a modular architecture (Huang et al., 2021; Wang et al., 2022; Shi et al., 2022; Huang et al., 2023; Shi et al., 2024), separating perception, prediction, planning, and control into independently optimized components. Recently, advances in large-scale multi-modal data and computational power have led to new paradigms, including end-to-end approaches (Hu et al., 2023; Jiang et al., 2023; Weng et al., 2024; Li et al., 2024; Tian et al., 2024) and Vision-Language-Action (VLA) models (Hwang et al., 2024; Xu et al., 2024; Zhou et al., 2025a; b), significantly enhancing system generalization and complex task performance. Despite these advancements, current systems still struggle with real-world diversity and replicating human driving preferences. Accordingly, within the VLA paradigm, comprehensive and rigorous evaluation of the entire pipeline is increasingly crucial in both academia and industry.
@@
Recent benchmarks and datasets represent significant progress, yet they still struggle to capture the diversity, complexity, and behavioral characteristics of real-world driving. Most existing benchmarks (Qian et al., 2024; Guo et al., 2024; Nie et al., 2024; Sima et al., 2024) are constructed from open-source datasets (Dosovitskiy et al., 2017; Caesar et al., 2020; Sun et al., 2020; Mao et al., 2021; Krasin et al., 2017; Shang et al., 2019), resulting in limited source variety. Typically, these datasets are designed for object-level perception tasks and thus overlook the contextual richness and human intent intrinsic to realistic driving decisions. In addition, critical and challenging scenarios—such as road merges and exits, interactions with pedestrians, and construction zones—remain largely underrepresented, making evaluation results less relevant to practical deployment risks. Furthermore, the distribution of driver behaviors is highly imbalanced, with simple maneuvers like going straight dominating the data, while more complex events are insufficiently covered, leading to inadequate assessment of challenging behaviors.
@@
Existing approaches to action ground truth annotation exhibit several limitations. Some works (Qian et al., 2024; Guo et al., 2024) do not provide action-level annotations and focus only on perception or understanding tasks. Other approaches (Kim et al., 2018; Malla et al., 2023; Sima et al., 2024; Xie et al., 2025) utilize manually annotated action labels, but such labels are often generated after driving behavior occurs and therefore do not faithfully reflect real-time driving intent and decisions. This gap in high-fidelity action labels restricts the reliability and realism of current evaluation.
@@
Regarding the design of evaluation systems, most existing benchmarks do not fully capture the core logic of driving decision-making. Some works focus primarily on isolated tasks such as video captioning (Kim et al., 2018), object recognition (Qian et al., 2024), or spatial understanding (Guo et al., 2024), without systematically covering the entire process from vision to action. For benchmarks that do address the full decision pipeline (Sima et al., 2024; Xie et al., 2025), a forward logic is often adopted, starting from perception and proceeding sequentially through prediction, planning, and action modules. However, this may not adequately represent a goal-driven paradigm that considers dependencies from the perspective of final decisions. As a result, current evaluation standards may not be closely aligned with realistic human driving decisions, highlighting the need for more comprehensive and goal-oriented evaluation frameworks.
@@
  • Driver-Contributed Broad-Coverage Driving Scenarios. DriveAction is constructed from real-world data proactively collected by drivers of autonomous vehicles, which fundamentally distinguishes it from existing benchmarks and provides a wide spectrum of both everyday and challenging driving scenarios. Manual curation guarantees a comprehensive and representative collection of driving scenarios and actions.
  • Human Driving Preference-Aligned Ground Truth. Action labels are collected directly from real-time driving operations, faithfully capturing human intent at the moment of decision-making. To align with the output granularity of end-to-end large models, these labels are discretized into high-level actions, which reflect the categorical nature of human driving decisions. All labels are manually verified to ensure validity, with erroneous, unreasonable, or illegal behaviors excluded.
@@ 2.1 Autonomous Driving Models
Autonomous driving was initially approached through modular systems (Huang et al., 2021; Wang et al., 2022; Shi et al., 2022; Huang et al., 2023; Shi et al., 2024), in which perception, prediction, planning, and control modules were developed and optimized independently. Subsequently, research progressed towards end-to-end architectures (Hu et al., 2023; Jiang et al., 2023; Weng et al., 2024; Li et al., 2024) that directly learn mappings from raw sensory inputs to actions. With advances in vision-language models (VLMs), hybrid approaches have emerged, in which VLMs are integrated into end-to-end architectures as independent modules that provide low-frequency driving suggestions (Tian et al., 2024). Most recently, the VLA paradigm (Hwang et al., 2024; Xu et al., 2024; Zhou et al., 2025a; b) has been established, enabling deeper integration of end-to-end models and VLMs for richer context understanding and greater generalization, further highlighting the growing need for dedicated and comprehensive evaluation benchmarks.
@@ 2.2 Language-Related Benchmarks for Autonomous Driving
Existing works can be grouped into four categories. The first focuses on video captioning and explanation, represented by BDD-X (Kim et al., 2018), which links actions with textual descriptions, and DRAMA (Malla et al., 2023), which uses question-answer annotations to identify risk objects and their corresponding causes. The second emphasizes spatial perception. For example, NuScenesQA (Qian et al., 2024) is designed for object-level and multi-modal question answering, while DriveMLLM (Guo et al., 2024) evaluates spatial and localization capabilities. The third focuses on general scene understanding, including TextVQA (Singh et al., 2019), Next-qa (Xiao et al., 2021), and RealWorldQA (xAI, 2024), which evaluate models on text recognition, video behavior comprehension, and visual understanding in real-world scenarios. The fourth emphasizes evaluation across the entire autonomous driving pipeline, including Reason2Drive (Nie et al., 2024), DriveLM (Sima et al., 2024), and DriveBench (Xie et al., 2025), which use a forward logic starting from perception, thus insufficiently aligned with goal-driven dependencies and realistic human driving behaviors.
@@
BDD-X Caption/explanation Self-collected Manual None
DRAMA Caption/explanation Self-collected Manual Chain
NuScenesQA spatial perception nuScenes None None
DriveMLLM spatial perception nuScenes None None
TextVQA general understanding Open Images None None
Next-qa general understanding VidOR None None
RealWorldQA general understanding Self-collected None None
Reason2Drive AD System nuScenes, Waymo, ONCE Open source Chain
DriveLM AD System nuScenes, CARLA Manual Graph
DriveBench AD System DriveLM Manual Graph
DriveAction AD System Driver-contributed Real-time operations Tree
@@ 3 DriveAction
Inspired by existing benchmarks, we introduce DriveAction, the first action-driven benchmark specifically designed for VLA models, leveraging real-world driving preferences. The following sections provide a comprehensive description of DriveAction. Specifically, Section 3.1 highlights how driver-contributed and carefully curated data enable extensive scenario coverage and diverse action representation. Section 3.2 details action labels aligned with human-like decision making, collected from real-time driving operations and validated through multi-stage review. Section 3.3 introduces the action-rooted, tree-structured evaluation framework, supporting flexible and rigorous assessment of models across the full V-L-A pipeline.
@@ 3.1 Driver-Contributed Broad-Coverage Driving Scenarios
As shown in the Source column of Table 1, DriveAction uniquely aggregates real-world data collected by drivers of company-operated autonomous vehicles, in contrast to previous benchmarks that rely on self-collected or open-source data. Our dataset covers 148 cities and includes records of the complete lineup of mass-produced vehicles in our deployment. To ensure a comprehensive and representative collection of driving scenarios and actions, we perform multiple rounds of manual selection and quality control.
Table 2 summarizes the seven key scenario categories in DriveAction, each paired with representative actions and concise descriptions to illustrate the coverage and annotation strategy. These scenarios span a wide range of real-world driving conditions, including ramp and side road merging/splitting, as well as navigation- and efficiency-driven lane changes. The coverage extends beyond typical urban and highway environments to encompass challenging contexts such as complex intersections, construction zones, congestion, and interactions with vulnerable road users. Each scenario is associated with a variety of fine-grained actions—such as lane changes, deceleration, and bypass maneuvers—enabling detailed analysis of decision-making across diverse driving situations.
@@
Bypass VRU Bypass left/right For bypassing vulnerable road users
@@ 3.2 Human Driving Preference-Aligned Ground Truth
DriveAction derives action labels directly from actual driving operations, enabling accurate, real-time capture of driver intent and decisions. This contrasts with previous benchmarks, which rely on manual or open-source annotation, as indicated in the Label column of Table 1.
DriveAction adopts discretized high-level actions as ground truth, which matches the output granularity of end-to-end large models and better reflects the categorical nature of human decisions. For example, in lane change scenarios, instead of relying on trajectories sampled at high frequency, DriveAction captures decisions made at key points—such as whether to initiate a lane change and which direction to take—better matching the lower decision frequency of large models. This design provides a more suitable and fair evaluation standard for current autonomous driving models.
@@ 3.3 Action-Rooted Tree-Structured Evaluation
To establish an evaluation framework that systematically evaluates models across the full spectrum of real-world driving decisions, we propose an action-rooted, tree-structured evaluation architecture in DriveAction, which dynamically maps complex driving actions to the required vision and language tasks. The integration of rich contextual information ensures that decisions are made within a complete and realistic environment, and the highly flexible evaluation system supports both comprehensive V-L-A evaluation and task-specific evaluation.
@@ 3.3.1 Task Definition
Figure 1 illustrates the task architecture of DriveAction, which uses the action space as the root, explicitly defines the corresponding language tasks required for each action, and further maps each language task to its dependent vision tasks. This hierarchy is organized into three layers: the top layer consists of action nodes, such as lane change and intersection turning, representing final decision outputs of the model. The middle layer consists of language tasks, such as navigation following and traffic light following, which provide scene understanding for each action. The bottom layer comprises vision tasks, responsible for detecting and recognizing key environmental elements, such as lanes, traffic signs, and obstacles. Although the evaluation framework is modeled as a dependency tree descending from actions, the underlying model inference still follows the conventional V-L-A order. This design offers a systematic and targeted evaluation structure, thus closely matching the information flow and reasoning processes of real-world autonomous driving systems.
DriveAction comprises 14 independent tasks, including 7 vision tasks and 7 language tasks, as illustrated in Figure 1. All tasks are evaluated in a question-answering (QA) format, including both selection and judgment modes. Figure 2 presents the QA distribution across vision, language, and action levels, ensuring sufficient and representative coverage within each layer of the VLA hierarchy. For instance, turn actions constitute a significant portion, which aligns with their high frequency in real-world scenarios. All task definitions and example QA pairs are detailed in the Appendix.
@@
Figure 3 uses the traffic sign task to exemplify the complete V-L-A evaluation pipeline in DriveAction. The process begins with detecting traffic signs, where the model is required to identify both their presence and type. Upon successful detection, the model must then interpret whether the sign is relevant to the current driving context and its own behavior. If the sign is deemed impactful, the model is expected to make an appropriate driving decision accordingly. Through this V-L-A pipeline, DriveAction comprehensively evaluates the model’s ability to transition from low-level perception to high-level understanding, and ultimately to informed decision-making.
@@ 3.3.2 Scenario Information Design
DriveAction incorporates key scenario context into each evaluation prompt, ensuring that models are assessed under complete and realistic conditions, and mitigating hallucinatory reasoning. Specifically, three types of scenario information are provided:
@@
  • Vehicle Speeds: Ego and target vehicle speeds, obtained from onboard sensors, quantify both the current and desired driving states. As dynamic information unavailable from images alone, they are essential for lane changes, overtaking, and acceleration decisions.
To assess the impact of scenario information, we examine the lane change task before an intersection under two conditions: with and without navigation instructions. As shown in Figure 4, the model reliably selects the correct lane when informed of an upcoming right turn, whereas lacking this guidance, it often makes ”hallucinated” decisions misaligned with the driving goal. This highlights the necessity of scenario information for reliable and context-aware evaluation in autonomous driving.